finding the center, foci, major and minor axis for the equation: 3x^2 + 9y^2 = 27

Answer by lwsshak3(11628)

finding the center, foci, major and minor axis for the equation:
3x^2 + 9y^2 = 27
divide by 27
x^2/9 + y^2/3 = 1
This is an equation of an ellipse with horizontal major axis.
Its standard form: (x-h)^2/a^2+(y-k)^2/b^2=1, a>b, (h,k)=(x,y) coordinates of center
center: (0,0)
a^2=9
a=3
length of major axis=2a=6
vertices: (0±a,0)=(±3,0)=(-3,0) and (3,0)
..
b^2=3
b=√3≈1.7
length of minor axis=2b=2√3
end points of minor axis: (0,±b)=(0,±1.7)=(0-1.7) and (0+1.7)
..
c^2=a^2-b^2=9-3=6
c=√6≈2.5
foci: (0±c,0)=(±2.5,0)=(-2.5,0) and (2.5,0)